If tanA1, tan A2, tanA3 and tanA4 are the roots of the equation
x^4 - x^3 sin2q + x^2 cos2q - x cosq - sinq = 0
then tan(A1 + A2 + A3 + A4) is equal to
(a)sinq
(B)cosq
(C)tanq
(D)cotq

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Ajay

let s1= sum of the roots taken one at a time = -(-sin2q)=sin2q

s2=sum of the roots taken two at a time = cos2q

s3=sum of the roots taken two at a time = -(-cosq)=cosq

s4= sum of the roots taken four at a time = -sinq

tan(A1+A2+A3+A4)= (s1-s3)/(1-s2-s4) = (sin2q-cosq)/(1-cos2q-sinq)

= (2sinq*cosq - cosq)/(2sin2q - sinq)

=cosq(2sinq-1)/sinq(2sinq-1)

=cotq

option D

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..